OpenStudy (anonymous):

Nobody can help me. This question is here like for 1 hour: Locus of points z: |z-2|-|z+2|>3

OpenStudy (experimentx):

x< -2

OpenStudy (anonymous):

how you got that?

OpenStudy (experimentx):

wait,

OpenStudy (experimentx):

it not 2, it's x<-3/2

OpenStudy (anonymous):

but explain how you got it

OpenStudy (experimentx):

hit and trial ,,, have to wait for theory

OpenStudy (anonymous):

i don't have correct answer. And not making any test. Just studying...

OpenStudy (experimentx):

|z-2|-|z+2| > |z|+|2|-|z+2| >|z|+|2|-|z|-|2| > 3 => 2|z| > 3 => z > 3/2 or -z < 3/2 case 1 doesn't work so case 2 ... hahah ... lol good night 3 am here

OpenStudy (anonymous):

I get something like this: \[28x ^{2} -14x +81>36y ^{2}\]

OpenStudy (experimentx):

have to sleep ... can we discuss tomorrow?

OpenStudy (anonymous):

|z-2|>|z|+|2|? kk, goodnight

OpenStudy (experimentx):

oops sorry mistake here => z > 3/2 (for positive z) or z < -3/2 (for negative z) good night

OpenStudy (anonymous):

|z-2|>|z|-|2|

OpenStudy (experimentx):

|z-2|>|z|+|2| wrong

OpenStudy (anonymous):

z is complex number

OpenStudy (experimentx):

Oo ,,, sorry .. wrong assumption. let's discuss tomorrow

OpenStudy (anonymous):

kk, thx

sam (.sam.):

|z-2|-|z+2|>3 |z-2|>3+|z+2| |x+yi-2|>3+|(x+yi)+2| |(x-2)^2+y^2|>3+|(x+2)^2+y^2| x^2-4x+4+y^2>3+(x^2+4x+4)+y^2 -4x+4>4x+7 -8x>3 x<3/8

OpenStudy (anonymous):

|(x-2)^2+y^2|>3+|(x+2)^2+y^2| ?? should't be square root?

OpenStudy (experimentx):

yeah sqrt is missing ..

OpenStudy (anonymous):

I can come to this expression: \[14x(x-1/2)>9(y^{2}-9/4)\] anybody recognize what's that??

sam (.sam.):

whoops forgot the square root \[\sqrt{(x-2)^2+y^2}>\sqrt{(x+2)^2+y^2}+3\]

OpenStudy (anonymous):

but then yout result is wrong

sam (.sam.):

you just need to sqrt both sides then the \[x<-\frac{3}{2} \]

OpenStudy (anonymous):

\[(x-2)^{2}+y ^{2}>(x+2)^{2}+y ^{2} +6\sqrt{(x+2)^{2}+y ^{2}} +9\]

OpenStudy (anonymous):

can you show me how you got that? @.Sam.

sam (.sam.):

which one?

OpenStudy (anonymous):

x<-3/2

OpenStudy (perl):

the one that was before the last ?